The sum of 3 numbers is 18. The sum of the first two numbers is 15. The first number is three times the size of the third number. What are the three numbers?
X+y+z=18

Found 2 solutions by Alan3354, Maths68:
Answer by Alan3354(69443)

The sum of 3 numbers is 18. The sum of the first two numbers is 15. The first number is three times the size of the third number. What are the three numbers?
X+y+z=18
x+y = 15
--> z = 3
---------
x = 3z
x = 9
----------
y = 6

Answer by Maths68(1474)

Let
Third Number = x
Second Number = y
First Number = 3x

The sum of 3 numbers is 18
x+3x+y=18
4x+y=18............(1)
The sum of the first two numbers is 15
3x+y=15..............(2)
Take (2)
3x+y=15
y=15-3x.........(3)
Put the value of y in (1)
4x+y=18
4x+(15-3x)=18
4x+15-3x=18
x=18-15
x=3
Put the value of x in (2)
3x+y=15
3(3)+y=15
9+y=15
y=15-9
y=6

Third Number = x = 3
Second Number = y = 6
First Number = 3x = 3(3) = 9